OpenStudy (anonymous):

vector question, BP,PC,AB,AC,AP are all vectors A, B, C are vertices of a triangle with a point P on the side of BC. Given that 3BP =4PC, find k if 3AB+ 4AC= kAP

OpenStudy (amistre64):

might help to draw out the triangle ....

OpenStudy (amistre64):



OpenStudy (amistre64):

AC = AB + BP + PC AP = AB + BP 3AB+ 4AB + 4BP + 4PC= kAB + kBP 3AB - kAB + 4AB = kBP - 4BP - 4PC 3AB - kAB + 4AB = kBP - 4BP - 3BP 7AB - kAB = kBP -7BP (7-k)AB = (k-7)BP those are just some ideas that come to mind ...